# Computing Wasserstein Barycenters through Gradient Flows

### v2

@@ 5 Conclusion
In this paper we have introduced a new framework for computing Wasserstein barycenters, based on Wasserstein gradient flows [ambrosio2008gradient, santambrogio2017euclidean]. Our approach addresses scalability issues in discrete barycenter solvers [cuturi2014fast], and is capable of regularizing the underlying barycenter through internal, potential, and interaction functionals. We presented gradient flow Algorithms 1 and 2, for empirical and Gaussian mixture measures, respectively. We further proved convergence guarantees based on the PL inequality.
We empirically tested our method against existing discrete [cuturi2014fast, montesuma2023multi] and neural-network [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing]-based barycenter solvers on computer vision, neuroscience, and chemical engineering benchmarks, demonstrating that our proposed methods consistently achieve state-of-the-art performance. Our empirical findings (Table 2) further demonstrate that incorporating label information in the OT objective is key for domain adaptation. This finding highlights an interesting gap in the literature, where neural net-based solvers are not capable of exploiting this information, and thus have sub-optimal performance in domain adaptation. We leave an investigation of this issue for future work.
Computing Wasserstein Barycenters through Gradient Flows
@@ Appendix A Introduction
Organization. We divide this supplementary material as follows. Section B provides a brief survey on the methods for computing Wasserstein barycenters. Section C.1 provides additional background on OT. Section C.2 covers additional details on the potential, internal, and interaction energies, as well as the barycenter objectives and their gradients. Section D provides missing proofs. Finally, section E and F provide additional experiments.
@@ B.1 Introduction
Given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, the Wasserstein barycenter problem was first introduced by [agueh2011barycenters] as the following minimization problem,
$P^{\star}=\underset{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr\}.$ (24)
This minimization problem is the analogous, in Wasserstein space, to the Euclidean barycenter problem $x\mapsto\sum_{k}\lambda_{k}d(x,x_{k})^{2}$, which is made possible since $(\mathcal{P}_{2,\text{ac}}(\Omega),\mathbb{W}_{2})$ is a metric space. Solving problem 24 is, in general, challenging. Except for the Gaussian case, there are no closed-form solutions. In the following, we review a few strategies for computing Wasserstein barycenters.
@@ B.2 Discrete Barycenter Solvers
$\hat{P}(z)=\sum_{i=1}^{n}a_{i}\delta(z-z_{i}^{(P)}).$ (25)
Here, $\{z_{i}^{(P)}\}_{i=1}^{n}$ is the support of $\hat{P}$ and $a=\{a_{i}:a_{i}\geq 0\text{, and }\sum a_{i}\}_{i=1}^{n}$ are the sample weights. There are two prevalent strategies when discretizing $P$. First, one can fix a grid/bins over the underlying space $\Omega$, and compute $a_{i}=P(z_{i}^{(P)})$ as the bin weight. This strategy is equivalent to estimating the density of $P$ over $\Omega$. Second, we can sample $z_{i}^{(P)}\overset{\text{i.i.d.}}{\sim}P$, in which case $a_{i}=\nicefrac{{1}}{{n}}$. In this paper, we deal primarily with free-support Wasserstein barycenters. Arguably, the first free-support algorithm was proposed by [cuturi2014fast]. Their strategy comes from plugging equation 25 into 24,
$Z^{(P),\star}=\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}\lVert z_{i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}\biggr\}.$ (26)
At this point, one should compare equation 26 with equation 17 in the main paper. Cuturi and Doucet [cuturi2014fast] propose solving this equation via a block coordinate descent strategy. Indeed, assume $\gamma_{k}=\text{OT}(P,Q_{k})$ has been calculated (e.g., through linear programming), then,
$\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}(z_{i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}.$
As a consequence, $\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=0$ leads to an update rule in terms of the barycentric projection of $P$ to $Q_{k}$, which is a discrete approximation of the Monge map [deb2021rates].
An important question comes when the samples from $P$ are feature-label joints. This is question was first raised by [montesuma2023multi], who proposed to use the ground-cost,
@@
The first paper to remark these computations was [montesuma2023multi], who proposed to couple these mappings,
$x_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(x_{i,\tau}^{(P)})\text{, and, }y_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)}).$ (28)
These iterations have a few nice properties with respect to the label term. First, the mappings $T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)})$ correspond to label propagation terms first defined in [redko2019optimal]. Second, as a consequence of the first point, the iterations respect the constraint $y_{i,\tau}^{(P)}\in\Delta_{n_{\text{classes}}}$. Third, these iterations can be understood as fixed-point iterations in the sense of [alvarez2016fixed]. The authors use this algorithm as the building block of their dictionary learning strategy.
The main drawback of the iterations in equation 28 is when the input measures have imbalanced classes. As we show in the additional experiments in Section E.2, OT, due the mass preservation constraints, is forced to match samples from different classes, leading to deformed barycenters with fuzzy labels (see Figure 8).
@@ B.3 Neural Barycenter Solvers
In this section, we cover a collection of methods that use neural networks as the foundation of computing Wasserstein barycenters. These methods are oftentimes referred to as ”continuous”, thus relying on the idea that the calculated barycenter is a continuous measure. Note that this is different than claiming that these barycenters solve a continuous problem. In fact, most methods are formulated in terms of expectations of the involved measures, which are approximated through mini-batching.
Generative Modeling. One of the first methods was proposed by [cohen2020estimating], who proposed viewing equation 24 through the lens of generative modeling. Let $g_{\theta}$ denote a neural network with parameters $\theta$. Then,
@@
where $P_{\theta}=g_{\theta,\sharp}P_{0}$ is the push-forward of a latent measure $P_{0}$ by the generator $g_{\theta}$. Therefore, the Wasserstein barycenter is parametrized via the weights of $g_{\theta}$. In comparison, methods such as [cuturi2014fast] and [montesuma2023multi] parametrize the barycenter through its samples. The underlying idea is sampling from each $Q_{k}$ and $P_{0}=\mathcal{N}(0,\sigma^{2}I)$.
Input Convex Neural Nets. More refined settings exploit the properties of the 2-Wasserstein distance with the squared Euclidean ground-cost. Indeed, note that,
@@
where $(f,g)\in\Phi=\{f(x)+g(y)\leq\lVert x-y\rVert_{2}^{2}\}$ are called potentials. The main insight of [makkuva2020optimal] is rewritting the minimization problem in the r.h.s. of equation 29 as,
@@
Based on these equations, [fan2020scalable] propose solving,
@@
An important aspect of equation 30 is the minimization with respect convex functions $f_{k}$ and $g_{k}$. Indeed, these are the potentials of the transportation problem between the barycenter $P$ and each $Q_{k}$. More specifically, In light of Brenier’s theorem [brenier1991polar], $\nabla g_{k}$ is the Monge map from $Q_{k}$ to the barycenter $P$. These elements motivate [fan2020scalable] using ICNNs, as in [makkuva2020optimal], for the optimization problem in equation 30.
Continuous 2-Wasserstein Barycenter. We now cover a family of methods relying on the fixed-point view of Wasserstein barycenters introduced by [alvarez2016fixed]. Indeed, these authors show that the Wasserstein barycenter is a fixed point of the operator $\Psi:P\mapsto(\sum_{k}\lambda_{k}T_{P\rightarrow Q_{k}})_{\sharp}P$. This is actually the principle behind most of the discrete methods described in the previous section. Now, with respect to OT potentials, if $\Psi(P)=P$, then one has,
$\sum_{k=1}^{K}\lambda_{k}T_{Q_{k}\rightarrow P}^{\star}(z)=x\overset{(a)}{\implies}\sum_{k=1}^{K}\lambda_{k}\nabla g_{k}(z)=z\overset{(b)}{\implies}\sum_{k=1}^{K}\lambda_{k}g_{k}(z)=\dfrac{\lVert z\rVert_{2}^{2}}{2}+c$
where implication (a) follows from Brenier’s theorem, and $g_{k}$ is the potential of the transportation problem from $Q_{k}$ to $P$, and implication (b) follows by integration. Here, since $c$ is an integration constant, one can conveniently take $c=0$. Now, denote $\text{Congruent}(\lambda)=\{(g_{1},\cdots,g_{K}):\sum_{k}\lambda_{k}g_{k}(z)=\nicefrac{{\lVert z\rVert_{2}^{2}}}{{2}}\}$ be the set of congruent potentials $(g_{1},\cdots,g_{K})$ with respect to coordinates $\lambda$. Using the dual formulation in equation 29 and restricting the set of potentials to congruent potentials,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]+\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\biggr(\sum_{k=1}^{K}\lambda_{k}(\mathbb{E}_{z\sim Q_{k}}[f_{k}(z)]+\mathbb{E}_{z\sim P}[g_{k}])(z)\biggr).$
Now, taking the expectation on the congruence constraint, we have,
@@
which allowed [korotin2021continuous] to simplify $\mathbb{B}_{\mathcal{Q}}(P)$ into,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}[g_{k}(z)],$ (31)
which does not involve terms with respect to the barycentric measure $P$. While the approach of [korotin2021continuous] simplifies the barycenter objective to expectations with respect to the known measures $Q_{k}$, it introduces a challenging constraint – the congruence of potentials. This constraint is challenging, as it is a functional inequality over the whole space $\Omega$. Furthermore, it adds to the already constrained setting of fitting convex functions. [korotin2021continuous] proposes solving it approximately, by penalizing potentials that do not meet the congruence condition. There is a circular dependence here, as one needs an estimate of the Wasserstein barycenter to estimate the constraint violation.
Neural (Weak) Optimal Transport. Weak OT [gozlan2017kantorovich, backhoff2019existence] comes from the interpretation of the Kantorovich formulation as,
@@
this formulation allows for more general costs $C:\Omega\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$, i.e., costs that take a sample and a probability measure as inputs. In this sense, [kolesov2024estimating] extends equation 31 to,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma(\cdot|z)}(g_{k}(z^{\prime}))]\biggr),$ (33)
which holds for congruent potentials $\sum_{k}\lambda_{k}g_{k}=0$. As with equation 31, $\mathbb{B}_{\mathcal{Q}}$ does not depend explicitly on the unknown barycentric measure $P$. As such, the authors compute the barycenter through a max-min optimization problem,
$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma_{k}(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}(g_{k}(z^{\prime}))]\biggr)\biggr\}.$
@@
where $C_{\text{classical}}$ allows us to retrieve the original OT problem. $C_{\text{KL}}$ is an alternative, when modeling $\gamma(\cdot|z)=\mathcal{N}(\mu(z),\sigma(z))$. On top of this reformulation, the authors in [kolesov2024estimating] propose approximating $\gamma(\cdot|z)$ through a mapping, as in [korotin2022neural], that is,
@@
where $\mathbb{S}$ is some prior measure, such as $\mathbb{S}=\mathcal{N}(0,\text{I})$. In this context, [gazdieva2024robust] extends this idea in the semi-unbalanced OT setting [liero2018optimal], where one of the mass preservation constraint of one of the marginals of $\gamma$ is relaxed. More specifically, given $\xi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$,
@@
where $P\ll Q$ means that $P$ is absolutely continuous with respect to $Q$. We denote $\bar{\xi}(z^{\prime})=\text{sup}_{z\in\mathbb{R}}\{zz^{\prime}-\xi(z)\}$ to be its convex conjugate. With that in mind, [gazdieva2024robust] shows, in Theorem 1 and Corollaries 1 and 2, that minimizing,
$\underset{P\in\mathcal{P}_{2,\text{ac}}}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\biggr(\int_{\Omega}\int_{\Omega}c(z,z^{\prime})d\gamma(z,z^{\prime})+D_{\xi}(\gamma(\cdot|z)||Q_{k})\biggr),$
@@
$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}\biggr[-\bar{\xi}_{k}\biggr(-\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}[c_{k}(z,z^{\prime})-g_{k}(z^{\prime})]\biggr)\biggr]\biggr)\biggr\}.$
Normalizing Flows. This approach was proposed by [visentin2025computing], and relies on the analysis of [gangbo1996geometry]. Their main idea is re-writing the Wasserstein distance in terms of functions $f,g$,
$\mathbb{W}_{p}(P,Q)^{p}=\underset{f\in\text{Borel}(P_{0},P),g\in\text{Borel}(P_{0},Q)}{\text{inf}}\,\lVert f-g\rVert_{L_{p}(P_{0})}^{p},$ (34)

### v3

@@ 5 Conclusion
We have presented a new Wasserstein barycenter algorithm (Algorithm 1), inspired by the gradient flow literature santambrogio2015optimal, that addresses two persistent challenges: scalability and regularization. We have derived a new algorithm that is amenable to mini-batch OT fatras2021minibatch (Section 3.1). Under entropic OT, we further demonstrate that OT computations can be vectorized across the $K$ input measures, leading to further speedups (Figure 5).
Our framework is flexible, and accommodates modular plug-in regularization from internal, potential, and interaction energies (Section 3.3). Extensive experiments on domain adaptation demonstrate the value of our approach and that incorporating labels in the OT ground-cost is necessary for state-of-the-art performance (Table 2).
Finally, our work opens interesting research questions. First, our framework can be extended to more complex differentiable structures, such as Riemannian manifolds. Second, it opens the question on how to effectively integrate labels and other regularization schemes into neural network solvers.
@@ Appendix A Introduction
Organization. We divide this supplementary material as follows. Section B provides a brief survey on the methods for computing Wasserstein barycenters. Section C.1 provides additional background on OT. Section C.2 covers additional details on the potential, internal, and interaction energies, as well as the barycenter objectives and their gradients. Section D provides missing proofs. Finally, section F provide additional experiments.
@@ B.1 Introduction
Given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, the Wasserstein barycenter problem was first introduced by agueh2011barycenters as the following minimization problem,
$P^{\star}=\underset{P\in\mathcal{P}_{2}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{\epsilon}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2,\epsilon}(P,Q_{k})^{2}\biggr\},$ (25)
where $\epsilon\geq 0$ denotes the entropic regularization weight. This minimization problem is the analogous, in Wasserstein space, to the Euclidean barycenter problem $x\mapsto\sum_{k}\lambda_{k}d(x,x_{k})^{2}$, which is made possible since $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ is a metric space. Solving problem 25 is, in general, challenging. Except for the Gaussian case, there are no closed-form solutions. In the following, we review a few strategies for computing Wasserstein barycenters.
@@ B.2 Discrete Barycenter Solvers
$\hat{P}(z)=\sum_{i=1}^{n}a_{i}\delta(z-z_{i}^{(P)})\in\text{Emp}_{n}(\Omega).$ (26)
Here, $\{z_{i}^{(P)}\}_{i=1}^{n}$ is the support of $\hat{P}$ and $a=\{a_{i}:a_{i}\geq 0\text{, and }\sum a_{i}\}_{i=1}^{n}$ are the sample weights. There are two prevalent strategies when discretizing $P$. First, one can fix a grid/bins over the underlying space $\Omega$, and compute $a_{i}=P(z_{i}^{(P)})$ as the bin weight. This strategy is equivalent to estimating the density of $P$ over $\Omega$. Second, we can sample $z_{i}^{(P)}\overset{\text{i.i.d.}}{\sim}P$, in which case $a_{i}=\nicefrac{{1}}{{n}}$. In this paper, we deal primarily with free-support Wasserstein barycenters. Arguably, the first free-support algorithm was proposed by cuturi2014fast. Their strategy comes from plugging equation 26 into 25,
$Z^{(P),\star}$ $=\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{arg min}}\,\biggr\{\mathbb{B}_{\epsilon}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{\epsilon,k,i,j}\lVert z_{i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}\biggr\},$ (27)
subject to $\gamma_{\epsilon,k}^{\star}=\underset{\gamma\in\Gamma(P,Q_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{ij}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}+\epsilon\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{ij}\log\gamma_{ij}$
At this point, one should compare equation 27 with equation 18 in the main paper. cuturi2014fast propose solving this equation via a block coordinate descent strategy. Indeed, assume $\gamma_{k}=\text{OT}(P,Q_{k})$ has been calculated (e.g., through linear programming), then,
$\nabla_{z_{i}^{(P)}}\mathbb{B}_{\epsilon}(P|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}(z_{i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}.$
As a consequence, $\nabla_{z_{i}^{(P)}}\mathbb{B}_{\epsilon}(P|\mathcal{Q})=0$ leads to an update rule in terms of the barycentric projection of $P$ to $Q_{k}$, which is a discrete approximation of the Monge map deb2021rates.
An important question comes when the samples from $P$ are feature-label joints, first raised by montesuma2023multi, who proposed to use the ground-cost,
@@
The first paper to remark these computations was montesuma2023multi, who proposed to couple these mappings,
$x_{\tau+1,i}^{(P)}=\sum_{k=1}^{K}\lambda_{k}\biggr(\underbrace{n\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}x_{j}^{(Q_{k})}}_{T_{P\rightarrow Q_{k}}^{\star}(x_{\tau,i}^{(P)})}\biggr)\text{, and, }y_{\tau+1,i}^{(P)}=\sum_{k=1}^{K}\lambda_{k}\biggr(\underbrace{n\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}y_{j}^{(Q_{k})}}_{T_{P\rightarrow Q_{k}}^{\star}(y_{\tau,i}^{(P)})}\biggr).$ (29)
These iterations have a few nice properties with respect to the label term. First, the mappings $T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)})$ correspond to label propagation terms first defined in redko2019optimal. Second, since the summations in equation 29 are convex combinations of the labels, the update rules naturally suffice $y_{\tau,i}^{(P)}\in\Delta_{n_{\text{classes}}}$. Third, these iterations can be understood as fixed-point iterations in the sense of alvarez2016fixed. The authors use this algorithm as the building block of their dictionary learning strategy.
@@ B.3 Neural Barycenter Solvers
In this section, we cover a collection of methods that use neural networks as the foundation of computing Wasserstein barycenters. These methods are oftentimes referred to as "continuous", thus relying on the idea that the calculated barycenter is a continuous measure. Note that this is different than claiming that these barycenters solve a continuous problem. In fact, most methods are formulated in terms of expectations of the involved measures, which are approximated through mini-batching. Indeed, mini-batching is the main tool for scalability in Wasserstein barycenters.
Generative Modeling. One of the first methods was proposed by cohen2020estimating, who proposed viewing equation 25 through the lens of generative modeling. Let $g_{\theta}$ denote a neural network with parameters $\theta$. Then,
@@
where $P_{\theta}=g_{\theta,\sharp}P_{0}$ is the push-forward of a latent measure $P_{0}$ by the generator $g_{\theta}$. Therefore, the Wasserstein barycenter is parametrized via the weights of $g_{\theta}$. In comparison, methods such as cuturi2014fast and montesuma2023multi parametrize the barycenter through its samples. The underlying idea is sampling from each $Q_{k}$ and $P_{0}=\mathcal{N}(0,\sigma^{2}I)$.
Input Convex Neural Nets. More refined settings exploit the properties of the 2-Wasserstein distance with the squared Euclidean ground-cost. Indeed, from [makkuva2020optimal, Equation 4], exploiting the decomposition
$\lVert z-z^{\prime}\rVert_{2}^{2}=\lVert z\rVert_{2}^{2}+\lVert z^{\prime}\rVert_{2}^{2}-2\langle z,z^{\prime}\rangle,$
we can obtain,
@@
where $(f,g)\in\Phi=\{f(x)+g(y)\leq\lVert x-y\rVert_{2}^{2}\}$ are called potentials. The main insight of makkuva2020optimal is rewritting the minimization problem in the r.h.s. of equation 30 as,
@@
Based on these equations, fan2020scalable propose solving,
@@
An important aspect of equation 31 is the minimization with respect convex functions $f_{k}$ and $g_{k}$. Indeed, these are the potentials of the transportation problem between the barycenter $P$ and each $Q_{k}$. More specifically, In light of Brenier’s theorem brenier1991polar, $\nabla g_{k}$ is the Monge map from $Q_{k}$ to the barycenter $P$. These elements motivate fan2020scalable using Input Convex Neural Networks (ICNNs), as in makkuva2020optimal, for the optimization problem in equation 31.
Continuous 2-Wasserstein Barycenter. We now cover a family of methods relying on the fixed-point view of Wasserstein barycenters introduced by alvarez2016fixed. Indeed, these authors show that the Wasserstein barycenter is a fixed point of the operator $\Psi:P\mapsto(\sum_{k}\lambda_{k}T_{P\rightarrow Q_{k}})_{\sharp}P$. This is actually the principle behind most of the discrete methods described in the previous section. Now, with respect to OT potentials, if $\Psi(P)=P$, then one has,
$\sum_{k=1}^{K}\lambda_{k}T_{Q_{k}\rightarrow P}^{\star}(z)=z\overset{(a)}{\implies}\sum_{k=1}^{K}\lambda_{k}\nabla g_{k}(z)=z\overset{(b)}{\implies}\sum_{k=1}^{K}\lambda_{k}g_{k}(z)=\dfrac{\lVert z\rVert_{2}^{2}}{2}+c$
where implication (a) follows from Brenier’s theorem, and $g_{k}$ is the potential of the transportation problem from $Q_{k}$ to $P$, and implication (b) follows by integration. Here, since $c$ is an integration constant, one can conveniently take $c=0$. Now, denote $\text{Congruent}(\lambda)=\{g_{1:K}=(g_{1},\cdots,g_{K}):\sum_{k}\lambda_{k}g_{k}(z)=\nicefrac{{\lVert z\rVert_{2}^{2}}}{{2}}\}$ be the set of congruent potentials $g_{1:K}=(g_{1},\cdots,g_{K})$ with respect to coordinates $\lambda$. Using the dual formulation in equation 30 and restricting the set of potentials to congruent potentials,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]+\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{g_{1:K}\in\text{Congruent}(\lambda)}{\text{min}}\,\biggr(\sum_{k=1}^{K}\lambda_{k}(\mathbb{E}_{z\sim Q_{k}}[f_{k}(z)]+\mathbb{E}_{z\sim P}[g_{k}])(z)\biggr),$
where $g_{1:K}=(g_{1},\cdots,g_{K})$. Now, taking the expectation on the congruence constraint, we have,
@@
which allowed korotin2021continuous to simplify $\mathbb{B}(P|\mathcal{Q})$ into,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{g_{1:K}\in\text{Congruent}(\lambda)}{\text{min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}[g_{k}(z)],$ (32)
which does not involve terms with respect to the barycentric measure $P$. While the approach of korotin2021continuous simplifies the barycenter objective to expectations with respect to the known measures $Q_{k}$, it introduces a challenging constraint – the congruence of potentials. This constraint is challenging, as it is a functional inequality over the whole space $\Omega$. Furthermore, it adds to the already constrained setting of fitting convex functions. korotin2021continuous proposes solving it approximately, by penalizing potentials that do not meet the congruence condition. There is a circular dependence here, as one needs an estimate of the Wasserstein barycenter to estimate the constraint violation.
Neural (Weak) Optimal Transport. Weak OT gozlan2017kantorovich, backhoff2019existence comes from the interpretation of the Kantorovich formulation as,
@@
this formulation allows for more general costs $C:\Omega\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$, i.e., costs that take a sample and a probability measure as inputs. In this sense, kolesov2024estimating extends equation 32 to,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma(\cdot|z)}(g_{k}(z^{\prime}))]\biggr),$ (34)
which holds for congruent potentials $\sum_{k}\lambda_{k}g_{k}=0$. As with equation 32, $\mathbb{B}$ does not depend explicitly on the unknown barycentric measure $P$. As such, the authors compute the barycenter through a max-min optimization problem,
$\underset{\{g_{1:K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma_{k}(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}(g_{k}(z^{\prime}))]\biggr)\biggr\}.$
@@
where $C_{\text{classical}}$ allows us to retrieve the original OT problem. $C_{\text{KL}}$ is an alternative, when modeling $\gamma(\cdot|z)=\mathcal{N}(\mu(z),\sigma(z))$. On top of this reformulation, the authors in kolesov2024estimating propose approximating $\gamma(\cdot|z)$ through a mapping, as in korotin2022neural, that is,
@@
where $\mathbb{S}$ is some prior measure, such as $\mathbb{S}=\mathcal{N}(0,\text{I})$. In this context, gazdieva2024robust extends this idea in the semi-unbalanced OT setting liero2018optimal, where one of the mass preservation constraint of one of the marginals of $\gamma$ is relaxed. More specifically, given $\xi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$,
@@
where $P\ll Q$ means that $P$ is absolutely continuous with respect to $Q$. We denote $\bar{\xi}(z^{\prime})=\text{sup}_{z\in\mathbb{R}}\{zz^{\prime}-\xi(z)\}$ to be its convex conjugate. With that in mind, gazdieva2024robust shows, in Theorem 1 and Corollaries 1 and 2, that minimizing,
$\underset{P\in\mathcal{P}_{2}}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\biggr(\int_{\Omega}\int_{\Omega}c(z,z^{\prime})d\gamma(z,z^{\prime})+D_{\xi}(\gamma(\cdot|z)||Q_{k})\biggr),$
@@
$\underset{\{g_{1:K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}\biggr[-\bar{\xi}_{k}\biggr(-\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}[c_{k}(z,z^{\prime})-g_{k}(z^{\prime})]\biggr)\biggr]\biggr)\biggr\}.$
Normalizing Flows. This approach was proposed by visentin2025computing, and relies on the analysis of gangbo1996geometry. Their main idea is re-writing the Wasserstein distance in terms of functions $f,g$,
$\mathbb{W}_{2}(P,Q)^{2}=\underset{f\in\text{Borel}(P_{0},P),g\in\text{Borel}(P_{0},Q)}{\text{inf}}\,\lVert f-g\rVert_{L_{2}(P_{0})}^{2},$ (35)